OpenStudy (anonymous):

circumference of a circle solve for r c=2 pie r

OpenStudy (anonymous):

\[c = 2\pi r\]\[\implies \frac{1}{2\pi}\cdot c = \frac{1}{2\pi}\cdot 2\pi r\]\[\implies r = \frac{c}{2\pi}\]

OpenStudy (anonymous):

thanks